# Spatial averaging for light reflection and transmission through cold atom arrays

### v1

@@ Abstract
We theoretically and computationally investigate the role that the spatial spread of atoms plays in the transmission and reflection of weak light from atom arrays. In particular, we investigate whether coherent wave functions for the atoms’ positions leads to different results from a thermal distribution with the same spatial spread. We find that the coherence is not relevant when the light is weak and the electronic states evolve on time scales shorter than the oscillation period of the atoms in their traps. Full numerical calculations and derivations using the sudden approximation show that reflection and transmission agree with the simple averaging over atom positions. For parameters outside these restrictions, the simple spatial averaging my lead to inaccurate results.
@@ I Introduction
Inspired by advances, several groups have investigated scenarios involving light interacting with many atoms in regular arrangements. Chang et al. (2004, 2012); Jenkins and Ruostekoski (2012); Meir et al. (2014); Bettles et al. (2016); Sutherland and Robicheaux (2016); Facchinetti et al. (2016); Shahmoon et al. (2017); Asenjo-Garcia et al. (2017); Perczel et al. (2017a, b); Bettles et al. (2017); Grankin et al. (2018); Chang et al. (2018); Manzoni et al. (2018); Asenjo-Garcia et al. (2019); Henriet et al. (2019); Needham et al. (2019); Qu and Rey (2019); Zhang and Mølmer (2019); Masson et al. (2020); Javanainen and Rajapakse (2019); Guimond et al. (2019); Moreno-Cardoner et al. (2019); Williamson et al. (2020); Williamson and Ruostekoski (2020); Bettles et al. (2020); Ballantine and Ruostekoski (2020); Bekenstein et al. (2020); Rui et al. (2020); Cidrim et al. (2020); Masson and Asenjo-Garcia (2020); Ballantine and Ruostekoski (2021a); Cano (2021); Patti et al. (2021); Robicheaux (2021); Ballantine and Ruostekoski (2021b); Rubies-Bigorda and Yelin (2022); Masson and Asenjo-Garcia (2022); Sierra et al. (2022); Buckley-Bonanno et al. (2022); Ballantine and Ruostekoski (2022); Rubies-Bigorda et al. (2022); Moreno-Cardoner et al. (2022); Mann and Mariani (2022); Gutiérrez-Jáuregui and Asenjo-Garcia (2022a); Fernández-Fernández and González-Tudela (2022); Walther et al. (2022); de Paz et al. (2022); Gutiérrez-Jáuregui and Asenjo-Garcia (2022b); Ribeiro and Gardiner (2022); Wang et al. (2022); Pedersen et al. (2023); Rubies-Bigorda et al. (2023); Ruostekoski (2023); Sheremet et al. (2023); De Paz and Huidobro (2023); Fayard et al. (2023); Cech et al. (2023); Solomons and Shahmoon (2023); Baßler et al. (2023); Han et al. (2023); Srakaew et al. (2023); Scheil et al. (2023); Ostermann et al. (2024); Solomons et al. (2024); Wang et al. (2024); Ben-Maimon et al. (2024); Shah et al. (2024) Exciting possibilities arise when the separation of atoms is less than a wavelength of the light. Most of the theoretical investigations assume the atoms are on a perfect grid with zero spatial deviation which is not possible: the wave function for the center of mass motion must at least have a spread from the ground state wave function. For atoms not cooled to the motional ground state, the spread in positions is larger. Some of the calculations treated spatial deviations by fixing the atoms in space with random positions given by the spatial distribution function and averaging over the possible positions. While this prescription is reasonable, there has not been a clear derivation of how the spread of positions should be treated, especially in the limit the atoms are cooled to the motional ground state. Does the coherence of the vibrational wave function require special treatment? At a technical level, are the results the same for two cases where the spatial part of the density matrix has diagonal $\rho(x,x)$ the same but off-diagonal $\rho(x,x^{\prime})$ substantially different? The purpose of this paper is to provide a clear derivation and numerical examples of how to treat the spatial spread of atoms in arrays. The examples we give will be for when the light and atoms reach steady state although this condition is not necessary.
@@
For a general treatment, one of the conditions needs to be that the atoms scatter or reflect few photons while they evolve into the electronic steady state. If this condition is not satisfied, the wave function or density matrix associated with the atoms’ positions will evolve. This clearly leads to results that don’t solely depend on the positions of the atoms fixed in space. Typically, the scattering or reflection leads to heating the spatial degrees of freedom although special detuning and geometry could lead to cooling. This condition can be satisfied by decreasing the intensity of the incident light. Therefore, except for very subradiant cases, the one atom Rabi frequency will be much less than the one atom decay rate: $\Omega\ll\Gamma$.
@@
The treatment below will focus on the case where the trap frequency, $\omega_{t}$, is much less than the decay rate of the system. Except for very subradiant cases, this leads to $\omega_{t}\ll\Gamma$. In typical atom traps, the trap frequency in the plane of the atom array is different from that perpendicular to the plane of the atom array. Often, the in-plane frequency is larger than that perpendicular to the plane leading to a direction dependence to $\omega_{t}$. For the condition $\omega_{t}\ll\Gamma$, the largest of the frequencies should be used. Most experimental arrangements will satisfy this condition. We also make this restriction because smaller $\omega_{t}$ leads to larger spatial spreads which are more crucial to treat correctly.
@@
We show that within these conditions the proper treatment of the light plus atom system is to calculate the light properties for atoms fixed in space and then average the light properties over the positions of the atoms weighted by their position distribution. We used the sudden approximation to derive this result, Sec. III. We also numerically computed the reflection and transmission in simple cases using a full density matrix treatment, Sec. IV. The treatment is the same whether the spread in positions is from being in a thermal state or being in one vibrational eigenstate. Restricting the calculation to the ground vibrational state is a different approximation that can lead to qualitatively wrong changes in the properties of the light (transmission, reflection, etc.). For example, we describe a case in Sec. IV where even if the recoil energy is 3,000 times smaller than the vibrational energy spacing, performing a density matrix calculation with all atoms restricted to the ground state leads to qualitatively wrong changes in the light properties.
@@ II.2 Transmitted and reflected light
Since we are interested in the effect of atoms’ positions on the interaction of light, we also need to have a form for the operators that need to be evaluated to compute the intensity of light at a position away from the atoms. The transmitted and reflected light in steady state is somewhat undefined because for most cases the interaction with the light will lead to the atoms’ motion heating with time. Strictly speaking, there is no steady state. However, in most experiments, it is assumed that the light is weak enough that the internal states of the atom reach steady state before vibrations appreciable change. Thus, we will investigate the reflection and transmission on this time scale by choosing the Rabi frequency to be much smaller than $\Gamma$.
@@ III Sudden Approximation
with $\psi_{n_{j}}(\bm{r}_{j})$ the vibrational wave function for atom $j$. At this point, there are no approximations. This form is much more difficult to use in a fully numerical treatment because many more points in real space, $\bm{r}_{1},\bm{r}_{2}...$, are required for converged calculations compared to the vibrational states of Eq. (13). The trapping Hamiltonian in Eq. (4) leads to coherences between $\bm{r}_{j}$ and $\bm{r}^{\prime}_{j}$ and the transfer of amplitude between different points in the $3N$ space. To be clear, from completeness arguments, this formulation must give the same result as using the vibrational basis functions, Eq. (13).
@@
Although this result is simple, it does lead to somewhat surprising conclusions. For example, suppose all atoms start in the vibrational ground state and the parameter $\eta=k_{0}\sqrt{\hbar/(2M\omega_{t})}$ is tiny. Since the recoil energy over the vibrational spacing is $\eta^{2}$, one might expect that one could use the approximation that the atoms are in the ground state throughout the evolution. As will be seen in the next section, this qualitatively misses the changes due to the spread in positions.
@@ IV Results
Changing the wavelength to $\lambda=2$ mm, leads to $\eta\simeq 0.056$ and $\sigma/\lambda\simeq 0.0126$. This gives a spread of positions that is 1.55 times larger than the calculations in Figs. 1 and 2 which will lead to a larger effect from averaging. The results are shown in Fig. 3 where it is clear that spatial averaging has a larger effect. In the plot are shown the points for calculations with $n_{max}=0$, 1, and 2 where it is clear that most of the change occurs when going from 0 to 1. In fact for both the $\lambda=3.1$ and 2 mm calculations, the $n_{max}=1$ results are good enough to reach better than 1% accuracy. For the 2 mm calculation at $\Delta=0.36\Gamma$, going from $n_{max}=0$ to 1 to 2 gives $1-R-T$ going from $2.4\times 10^{-2}$ to $2.3\times 10^{-3}$ to $3.7\times 10^{-4}$ and an error in $T$ going from $-8.0\times 10^{-2}$ to $6.4\times 10^{-3}$ to $-5.4\times 10^{-4}$.
Figure 3: Similar to Fig. 2 but for $\lambda=2$ mm which increases the relative size of the averaging region. The results of the $n_{max}=0$ (purple solid square for $T$, empty pink square for $R$), $n_{max}=1$ (orange solid circle for $T$, maroon empty circle for $R$), and $n_{max}=2$ (turquoise solid triangle for $T$, empty green triangle for $R$).

### v2

@@ Abstract
We theoretically and computationally investigate the role that the spatial spread of atoms plays in the transmission and reflection of weak light from atom arrays. In particular, we investigate whether coherent wave functions for the atoms’ positions leads to different results from a thermal distribution with the same spatial spread. We find that the coherence is not relevant when the light is weak and the electronic states evolve on time scales shorter than the oscillation period of the atoms in their traps. Full numerical calculations and derivations using the sudden approximation show that reflection and transmission agree with the simple averaging over atom positions for these conditions. For parameters outside these restrictions, the simple spatial averaging may lead to inaccurate results.
@@ I Introduction
Inspired by theoretical and experimental advances, several groups have investigated scenarios involving light interacting with many atoms in regular arrangements. Chang et al. (2004); Henriet et al. (2019); Qu and Rey (2019); Chang et al. (2012); Jenkins and Ruostekoski (2012); Sutherland and Robicheaux (2016); Facchinetti et al. (2016); Shahmoon et al. (2017); Perczel et al. (2017a, b); Bettles et al. (2017); Asenjo-Garcia et al. (2019); Needham et al. (2019); Zhang and Mølmer (2019); Masson et al. (2020); Javanainen and Rajapakse (2019); Moreno-Cardoner et al. (2019); Williamson and Ruostekoski (2020); Bettles et al. (2020); Ballantine and Ruostekoski (2020); Masson and Asenjo-Garcia (2020); Ballantine and Ruostekoski (2021a); Cano (2021); Robicheaux (2021); Ballantine and Ruostekoski (2021b); Rubies-Bigorda and Yelin (2022); Masson and Asenjo-Garcia (2022); Sierra et al. (2022); Moreno-Cardoner et al. (2022); Mann and Mariani (2022); Gutiérrez-Jáuregui and Asenjo-Garcia (2022a); Fernández-Fernández and González-Tudela (2022); Walther et al. (2022); de Paz et al. (2022); Wang et al. (2022); Pedersen et al. (2023); De Paz and Huidobro (2023); Fayard et al. (2023); Solomons and Shahmoon (2023); Scheil et al. (2023); Ostermann et al. (2024); Wang et al. (2024); Ben-Maimon et al. (2024); Cech et al. (2023); Patti et al. (2021); Asenjo-Garcia et al. (2017); Manzoni et al. (2018); Guimond et al. (2019); Bekenstein et al. (2020); Cidrim et al. (2020); Buckley-Bonanno et al. (2022); Ballantine and Ruostekoski (2022); Rubies-Bigorda et al. (2022); Han et al. (2023); Shah et al. (2024); Meir et al. (2014); Grankin et al. (2018); Bettles et al. (2016); Chang et al. (2018); Williamson et al. (2020); Rui et al. (2020); Gutiérrez-Jáuregui and Asenjo-Garcia (2022b); Ribeiro and Gardiner (2022); Rubies-Bigorda et al. (2023); Baßler et al. (2023); Srakaew et al. (2023); Solomons et al. (2024); Ruostekoski (2023); Sheremet et al. (2023) Exciting possibilities arise when the separation of atoms is less than a wavelength of the light. Most of the theoretical investigations assume the atoms are on a perfect grid with zero spatial deviation.Chang et al. (2004, 2012); Jenkins and Ruostekoski (2012); Sutherland and Robicheaux (2016); Facchinetti et al. (2016); Shahmoon et al. (2017); Asenjo-Garcia et al. (2017); Perczel et al. (2017a, b); Bettles et al. (2017); Manzoni et al. (2018); Asenjo-Garcia et al. (2019); Henriet et al. (2019); Needham et al. (2019); Qu and Rey (2019); Zhang and Mølmer (2019); Masson et al. (2020); Javanainen and Rajapakse (2019); Guimond et al. (2019); Moreno-Cardoner et al. (2019); Williamson and Ruostekoski (2020); Bettles et al. (2020); Ballantine and Ruostekoski (2020); Bekenstein et al. (2020); Cidrim et al. (2020); Masson and Asenjo-Garcia (2020); Ballantine and Ruostekoski (2021a); Cano (2021); Patti et al. (2021); Robicheaux (2021); Ballantine and Ruostekoski (2021b); Rubies-Bigorda and Yelin (2022); Masson and Asenjo-Garcia (2022); Sierra et al. (2022); Buckley-Bonanno et al. (2022); Ballantine and Ruostekoski (2022); Rubies-Bigorda et al. (2022); Moreno-Cardoner et al. (2022); Mann and Mariani (2022); Gutiérrez-Jáuregui and Asenjo-Garcia (2022a); Fernández-Fernández and González-Tudela (2022); Walther et al. (2022); de Paz et al. (2022); Wang et al. (2022); Pedersen et al. (2023); De Paz and Huidobro (2023); Fayard et al. (2023); Cech et al. (2023); Solomons and Shahmoon (2023); Han et al. (2023); Scheil et al. (2023); Ostermann et al. (2024); Wang et al. (2024); Ben-Maimon et al. (2024); Shah et al. (2024) While this is useful for exploring basic concepts, it is not possible in practice: the wave function for the center of mass motion must at least have a spread from the ground state wave function. For atoms not cooled to the motional ground state, the spread in positions is larger and many treatments averaged over spatial distributions in addition to treating the perfect lattice.Meir et al. (2014); Bettles et al. (2016); Grankin et al. (2018); Chang et al. (2018); Williamson et al. (2020); Rui et al. (2020); Gutiérrez-Jáuregui and Asenjo-Garcia (2022b); Ribeiro and Gardiner (2022); Rubies-Bigorda et al. (2023); Ruostekoski (2023); Sheremet et al. (2023); Baßler et al. (2023); Srakaew et al. (2023); Solomons et al. (2024) There is a wide range of applications for arrays of atoms including clocksChang et al. (2004); Henriet et al. (2019); Qu and Rey (2019), mirrors, light manipulation, and excitations on a latticeChang et al. (2012); Jenkins and Ruostekoski (2012); Bettles et al. (2016); Sutherland and Robicheaux (2016); Facchinetti et al. (2016); Shahmoon et al. (2017); Perczel et al. (2017a, b); Bettles et al. (2017); Chang et al. (2018); Asenjo-Garcia et al. (2019); Needham et al. (2019); Zhang and Mølmer (2019); Masson et al. (2020); Javanainen and Rajapakse (2019); Moreno-Cardoner et al. (2019); Williamson et al. (2020); Williamson and Ruostekoski (2020); Bettles et al. (2020); Ballantine and Ruostekoski (2020); Rui et al. (2020); Masson and Asenjo-Garcia (2020); Ballantine and Ruostekoski (2021a); Cano (2021); Robicheaux (2021); Ballantine and Ruostekoski (2021b); Rubies-Bigorda and Yelin (2022); Masson and Asenjo-Garcia (2022); Sierra et al. (2022); Moreno-Cardoner et al. (2022); Mann and Mariani (2022); Gutiérrez-Jáuregui and Asenjo-Garcia (2022a); Fernández-Fernández and González-Tudela (2022); Walther et al. (2022); de Paz et al. (2022); Gutiérrez-Jáuregui and Asenjo-Garcia (2022b); Ribeiro and Gardiner (2022); Wang et al. (2022); Pedersen et al. (2023); Rubies-Bigorda et al. (2023); De Paz and Huidobro (2023); Fayard et al. (2023); Solomons and Shahmoon (2023); Baßler et al. (2023); Srakaew et al. (2023); Scheil et al. (2023); Ostermann et al. (2024); Solomons et al. (2024); Wang et al. (2024); Ben-Maimon et al. (2024), collective Lamb shiftMeir et al. (2014), and quantum information applicationsAsenjo-Garcia et al. (2017); Grankin et al. (2018); Manzoni et al. (2018); Guimond et al. (2019); Bekenstein et al. (2020); Cidrim et al. (2020); Patti et al. (2021); Buckley-Bonanno et al. (2022); Ballantine and Ruostekoski (2022); Rubies-Bigorda et al. (2022); Cech et al. (2023); Han et al. (2023); Shah et al. (2024). Recent overviews cover important aspects of atom arrays.Ruostekoski (2023); Sheremet et al. (2023)
Some of the calculations treated spatial deviations by fixing the atoms in space with random positions given by the spatial distribution function and averaging over the possible positions. While this prescription is reasonable, there has not been a clear derivation of how the spread of positions should be treated, especially in the limit the atoms are cooled to the motional ground state. Does the coherence of the vibrational wave function require special treatment? At a technical level, are the results the same for two cases where the spatial part of the density matrix has diagonal $\rho(x,x)$ the same but off-diagonal $\rho(x,x^{\prime})$ substantially different? The purpose of this paper is to provide a clear derivation and numerical examples of how to treat the spatial spread of atoms in arrays. The examples we give will be for when the light and atoms reach steady state although this condition is not necessary.
@@
For a general treatment, one of the conditions needs to be that the atoms scatter or reflect few photons while they evolve into the electronic steady state. If this condition is not satisfied, the wave function or density matrix associated with the atoms’ positions will evolve. This clearly leads to results that don’t solely depend on the positions of the atoms fixed in space. Typically, the scattering or reflection leads to heating the spatial degrees of freedom although special detuning and geometry could lead to cooling. This condition can be satisfied by decreasing the intensity of the incident light. Therefore, except for very subradiant cases, the approximation will work well when the one atom Rabi frequency is much less than the one atom decay rate: $\Omega\ll\Gamma$.
@@
The treatment below will focus on the case where the trap frequency, $\omega_{t}$, is much less than the decay rate of the system. Except for very subradiant cases, this leads to $\omega_{t}\ll\Gamma$. In typical atom traps, the trap frequency in the plane of the atom array is different from that perpendicular to the plane of the atom array. Often, the in-plane frequency is larger than that perpendicular to the plane, leading to a direction dependence to $\omega_{t}$. For the condition $\omega_{t}\ll\Gamma$, the largest of the frequencies should be used. Most experimental arrangements will satisfy this condition. We also make this restriction because smaller $\omega_{t}$ leads to larger spatial spreads which are more crucial to treat correctly.
@@
We show that within these conditions an accurate, approximate treatment of the light plus atom system is to calculate the light properties for atoms fixed in space and then average the light properties over the positions of the atoms weighted by their position distribution. We used the sudden approximation to derive this result, Sec. III, hence the condition that the atoms do not move substantially for the duration of the light-atom interaction. We also numerically computed the reflection and transmission in simple cases using a full density matrix treatment, Sec. IV, to illustrate the accuracy of the fixed in space approximation. The numerical treatment is the same whether the spread in positions is from being in a thermal state or being in one vibrational eigenstate. Restricting the density matrix calculation to the ground vibrational state is a different approximation that can lead to qualitatively wrong changes in the properties of the light (transmission, reflection, etc.). For example, we describe a case in Sec. IV where even if the recoil energy is 3,000 times smaller than the vibrational energy spacing, performing a density matrix calculation with all atoms restricted to the vibrational ground state leads to qualitatively wrong changes in the light properties.
@@ II.2 Transmitted and reflected light
Since we are interested in the effect of atoms’ positions on the interaction of light, we also need to have a form for the operators that can be evaluated to compute the intensity of light at a position away from the atoms. The transmitted and reflected light in steady state is somewhat undefined because for most cases the interaction with the light will lead to the atoms’ motion heating with time. Strictly speaking, there is no steady state. However, in most experiments, it is assumed that the light is weak enough that the internal states of the atom reach steady state before vibrations appreciable change. Thus, we will investigate the reflection and transmission on this time scale by choosing the Rabi frequency to be much smaller than $\Gamma$ in most of our examples.
@@ III Sudden Approximation
with $\psi_{n_{j}}(\bm{r}_{j})$ the vibrational wave function for atom $j$. At this point, there are no approximations. This form is much more difficult to use in a fully numerical treatment because many more points in real space, $\bm{r}_{1},\bm{r}_{2}...$, are required for converged calculations compared to the vibrational states of Eq. (13). The trapping Hamiltonian in Eq. (4) leads to coherences between $\bm{r}_{j}$ and $\bm{r}^{\prime}_{j}$ and the transfer of amplitude between different points in the $3N$ space. To be clear, from completeness arguments, this formulation must give the same result when converged as using the vibrational basis functions, Eq. (13).
@@
Although this result is simple, it does lead to somewhat surprising conclusions. For example, suppose all atoms start in the vibrational ground state and the parameter $\eta=k_{0}\sqrt{\hbar/(2M\omega_{t})}$ is tiny. Since the recoil energy over the vibrational spacing is $\eta^{2}$, one might expect that one could use the approximation that the atoms are in the ground state throughout the evolution. As will be seen in the Results section, this can qualitatively miss the changes due to the spread in positions.
III.1 Evolution time scale
The discussion of the sudden approximation depends on the duration of the light-atom interaction. Sometimes the situation involves the properties of the light once the atoms reach steady state in which case the duration is until the electronic states stop evolving. Sometimes the properties are required before steady state is reached which would shorten the duration. The accuracy of the sudden approximation requires the atoms to be fixed in space for this duration. By this we mean, the atom positions would not be expected to change in an important way on the time scale of the measurement of the light properties.
Weak light, $\Omega\ll\Gamma$, leads to the simplest case. For steady state properties, the electronic states will require several lifetimes. The relevant lifetime for light coupling to superradiant states is less than $1/\Gamma$ while light coupling to subradiant states leads to relevant lifetimes larger than $1/\Gamma$. For this case, the condition will often be $(10-100)/\Gamma\ll 1/\omega_{t}$ depending on the extent of subradiance.
If the light is intense and/or the $\eta$ becomes large, the recoil of the atoms or dipole-dipole forces can also become important. The recoil or the dipole-dipole interactions could cause the atoms to accelerate so that the atom velocity increases substantially. The atom separation divided by the increased speed then becomes the comparison time scale, not $1/\omega_{t}$.
Lastly, we restrict the atoms to be cold. If, the atoms could be initially moving quickly, i.e. higher temperature, then the electronic states need to evolve on time scales fast compared the atom separation divided by the speed of the atoms. For this case, Doppler effects could lead to the collective light-atom interaction being relatively uninteresting.
@@ IV Results
Changing the wavelength to $\lambda=2$ mm, leads to $\eta\simeq 0.056$ and $\sigma/\lambda\simeq 0.0126$. This gives a spread of positions that is 1.55 times larger than the calculations in Figs. 1 and 2 which will lead to a larger effect from averaging. The results are shown in Fig. 3 where it is clear that spatial averaging has a larger effect. In the plot are shown the points for calculations with $n_{max}=0$, 1, and 2 where it is clear that most of the change occurs when going from 0 to 1. In fact for both the $\lambda=3.1$ and 2 mm calculations, the $n_{max}=1$ results are good enough to reach better than 1% accuracy. For the 2 mm calculation at $\Delta=0.36\Gamma$, going from $n_{max}=0$ to 1 to 2 gives $1-R-T$ going from $2.4\times 10^{-2}$ to $2.3\times 10^{-3}$ to $3.7\times 10^{-4}$ and an error in $T$ going from $8.0\times 10^{-2}$ to $-6.4\times 10^{-3}$ to $5.4\times 10^{-4}$.
Figure 3: Similar to Fig. 2 but for $\lambda=2$ mm which increases the relative size of the averaging region. The red solid (blue long dash) line is the $R$ ($T$) probability for perfect separation while the black dotted (green dash) spatially average the positions. The results of the $n_{max}=0$ (purple solid square for $T$, empty pink square for $R$), $n_{max}=1$ (orange solid circle for $T$, maroon empty circle for $R$), and $n_{max}=2$ (turquoise solid triangle for $T$, empty green triangle for $R$).
We repeated the calculations that used $\lambda=2$ mm but with 3 atoms to ensure there was nothing special about the 2 atom case. Because there are more atoms, the separation does not need to be as close to an integer wavelength to get sensitivity to the atom positions. In Fig. 4, we show the transmission and reflection probability versus detuning for 3 atoms with average separation of $d=0.85\lambda$. The line and point types are the same as Fig. 3. As with the previous plot, these show the very strong error for the calculation with $n_{max}=0$ and progressive convergence toward the spatial averaging result with increasing $n_{max}$. At $\Delta=0.36\Gamma$, going from $n_{max}=0$ to 1 to 2 to 3 gives $1-R-T$ going from $6.3\times 10^{-2}$ to $3.1\times 10^{-3}$ to $7.7\times 10^{-4}$ to $1.3\times 10^{-4}$ and an error in $T$ going from $5.4\times 10^{-2}$ to $-5.3\times 10^{-3}$ to $2.0\times 10^{-4}$ to $-6.6\times 10^{-5}$.
Figure 4: Same as Fig. 3 but for 3 atoms with a separation of $d=0.85\lambda$.
@@
The sudden approximation also can be made to fail for larger values of the Rabi frequency, $\Omega$. We repeated the calculations for the parameters of Fig. 3 but with $\Omega=10^{-3}$, $0.01$, and $0.1\Gamma$. All of these cases have $\Omega\gg\omega_{t}$, and, for the fixed atom calculations, lead to steady state results for times much less than $1/\omega_{t}$. Except for the $\Omega=\Gamma/10$ case, the simple spatial averaging gave an excellent approximation to the converged vibrational calculation. However, the $\Omega=\Gamma/10$ case gave qualitatively different results between the simple spatial averaging and the converged vibrational calculation because the atoms substantially shift positions under the radiation pressure from the reflected and scattered photons.